40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Feb 14 17:40:27 2020, anchal, DailyProgress, ISS, Installed ISS on both paths using SR560s IMG_20200214_174127.jpgISS_Effect.pdf
    Reply  Sat Feb 15 18:57:43 2020, rana, DailyProgress, ISS, Installed ISS on both paths using SR560s 
       Reply  Thu Feb 20 17:42:23 2020, rana, DailyProgress, ISS, Installed New ISS on both paths using SR560s ISS_New_Analysis.pdf
          Reply  Sun Feb 23 18:19:32 2020, rana, DailyProgress, ISS, Installed New ISS on both paths using SR560s 
             Reply  Fri Feb 28 14:00:11 2020, anchal, DailyProgress, ISS, Installed New ISS on both paths using SR560s - PROBLEM 
                Reply  Sun Mar 1 16:34:31 2020, rana, DailyProgress, ISS, Installed New ISS on both paths using SR560s - PROBLEM 
                   Reply  Mon Mar 2 16:48:12 2020, anchal, DailyProgress, ISS, Follow-up, porbalem stopped happening, OLTF plot. NISS_OLTF.pdf
Message ID: 2538     Entry time: Fri Feb 14 17:40:27 2020     Reply to this: 2539
Author: anchal 
Type: DailyProgress 
Category: ISS 
Subject: Installed ISS on both paths using SR560s 

I've installed ISS on both paths using 3 SR560s each. Preliminary feedback is setup to get a stable loop. More optimization with TF analysis can be done further.


Path changes:

  • I replaced the unmarked waveplate in the north path right after the EOAM at (112, 39) with a zeroth-order quarter waveplate.
  • I adjusted these quarter waveplates at (112, 39) and (90, 24) to the half transmission through the PBS when EOAM inputs are shorted.
  • I followed the notes in Antonio's post (See CTN:1593 and CTN:1604).

ISS:

  • ISS is formed with three SR560s in series.
  • First SR560 is AC coupled, has a 2nd order zero at 1 Hz and gain of 100 at Low noise Mode.
  • The second SR560 is DC coupled and has a pole at 1 kHz (North) and 3 kHz (South) with gain 1 and High Dynamic Reserve setting.
  • The third SR560 is also DC coupled and has a pole at 300 kHz with gain 1 and High Dynamic Reserve setting.

Measurement:

  • I took the measurement of the spectrum from the output of transmission photodiodes.
  • I used the calibrated coupling factor I measured earlier to convert the measured voltage to watts of power right after the cavities. (See CTN:2528).
  • Measurement configuration files is attached.
  • We see about a factor of 10 improvement in the intensity noise between 100Hz to 1 kHz.
  • Note that the bad spikes notes in the beatnote spectrum at 480 Hz and around 900 Hz are present in the north transmission spectrum as well.
  • Effect on the beatnote noise would be evident after tonight's measurement at 3 am. In my preliminary measurements, I see some reduction in noise below 200 Hz but nothing much.

Data

Attachment 2: ISS_Effect.pdf  59 kB  | Hide | Hide all
ISS_Effect.pdf
ELOG V3.1.3-